OpenStudy (anonymous):

x^2+3x-4=0 solve medals will be awarded

OpenStudy (anonymous):

(x-1)(x+4)

OpenStudy (anonymous):

so x=1 and x=-4

OpenStudy (anonymous):

thanks guys.

OpenStudy (id21):

\[x^2+3x-4=0\] \[x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}=\frac{-3 \pm \sqrt {9+16}}{2}=1;-4\]